OpenStudy (anonymous):

any help is appreciated!! simplify 2-i/2+5i

OpenStudy (anonymous):

Hey, buddy.

OpenStudy (anonymous):

Okay, so there's this thing for complex numbers called the "complex conjugate." And that's a real fancy sounding name, but all it means is that we take the complex number, leave the real part the same, and change the sign of the imaginary part. Here's an example: 2+4i has the complex conjugate 2-4i 1-3i has the complex conjugate 1+3i -5 +i has the complex conjugate -5-i understand? Pretty simple stuff.

OpenStudy (anonymous):

ya looks good to me!...basically the flip flop of the original equation

OpenStudy (anonymous):

Yeah, you can think of it that way. So why do we care about the complex conjugate? Well, it's actually pretty cool. It's cool because any time we take a complex number and multiply it by its complex conjugate, the result is just a real number! Sweet, right? So I said that the complex conjugate of 2 +4i is 2-4i. Let's see what happens when we multiply those two things. I'm going to basically FOIL them. (2+4i)(2-4i) = 4 - 8i +8i -16i^2 = 4-16i^2 because the -8i and the +8i cancel =4-16(-1) because i^2 =-1 =4+16 =20 So we multiplied them and just got a regular ol' number. Sweet!

OpenStudy (anonymous):

lol ya i suppose that is pretty sweet!

OpenStudy (anonymous):

You're probably thinking, "Great, but what does that have to do with this problem?" Well, we have this fraction \[\frac{2-i}{2+5i}\] and we'd like to simplify it. The WAY we'll simplify it is by rewriting it so that it no longer has a complex number on the bottom. The complex conjugate gives us a really cool way to do that. We know that if we multiply by the same thing on top and bottom, it doesn't change the value of the fraction. ALSO, we know that if we multiply a complex number by its complex conjugate, we'll end up with just a regular number.

OpenStudy (anonymous):

So, you tell me what we're going to do.

OpenStudy (anonymous):

hmm...good question i suppose...multiply the complex number on the bottom by the top and bottom in order to cancel it out on the bottom and move it to the top and then solve for the top by using the complex conjugate?...im i kinda right or am i making up a whole new math lol

OpenStudy (anonymous):

am i*

OpenStudy (anonymous):

You might be mildly making a new math. Let's keep it simple. We have a complex number on the bottom. How do we make it into NOT a complex number?

OpenStudy (anonymous):

wave a magic wand lol.....im thinking cancel it out somehow..which i would normally do by multiplying the i by the top and bottom

OpenStudy (anonymous):

Did you forget about the complex conjugate?

OpenStudy (anonymous):

no...im just looking at some notes i got from my teacher...the complex conjugate still includes the i though...so how would that help if you were trying to get rid of it?...there's obviously a reason my user name is mathdummy haha....i tried to solve this early and i got (1/29)-(12/29)i...am i at least close??

OpenStudy (anonymous):

oo wait scratch that i see y you need the conjugate

OpenStudy (anonymous):

Well, maybe. Let's talk about the process though. That's the important part. How do we GET to the answer. Not WHAT IS the answer.

OpenStudy (anonymous):

multiple the conjugate of 2+5i by the top and bottom

OpenStudy (anonymous):

The key fact to understand here is that if we have a complex number and we want it to become just a real number we multiply it by its complex conjugate.

OpenStudy (anonymous):

Good, good. So we look at the bottom, say to ourselves, "Let's turn that into a real number by multiplying it by its conjugate." But THEN, we say to ourselves, "Oh wait, since I'm multiplying on bottom, I have to multiply on top to keep it the same."

OpenStudy (anonymous):

agreed..so when working it out you factor the denominator in this case (2+5i)(2-5i) and you get 4-10i+10i-25i...correct?

OpenStudy (anonymous):

and the 10i cancel out...so you end up with 4-(-25i)...or at least i did

OpenStudy (anonymous):

Tiny, but important mistake 4-10i+10i-25i^2 Then the -10i+10i cancel and you get just 4-25i^2

OpenStudy (anonymous):

oo right..oops..and in this case...in know the i^2=-1....but because its attached to the 25 can u still do that?....im assuming not...sadly...that would make my life so much easier haha

OpenStudy (anonymous):

You totally can do that! And you should =)

OpenStudy (anonymous):

oo yes! finally something about math that goes my way! haha...okay so i did that..and for my shiny new answer i got -1/29-12/29i...

OpenStudy (anonymous):

Okay, so I guess you already did the work for the top. Here it is, just to confirm it for you: (2-i)(2-5i) = 4 -10i-2i +5i^2 = 4-12i +5(-1) = 4 -12i -5 = -1-12i And then, since the bottom was 29, we have \[\frac{-1-12i}{29} = \frac{-1}{29}-\frac{12}{29}i\]

OpenStudy (anonymous):

yep that's what i got!..thank you so much haha if you aren't a teacher already then you should seriously consider becoming one

OpenStudy (anonymous):

Any time, my little pumpkin. =) And I am already a teacher! I subbed for this past year, and this year I'll have my own high school math classroom! c:

OpenStudy (anonymous):

well your sure ready for it! hahaha thanks again..im sure ill be in need of your assistance again soon:)